OpenStudy (anonymous):

What is the solution set of |2x – 3| – 8 = –1?

OpenStudy (anonymous):

Rearrange the equation to give: |2x-3| = -1+8 = 7 What values of |2x-3| = 7? As before, you are looking for alues of x which make 2x-3 = -7 and 2x-3 = +7 Why are both +7 and =7 used? The | | means absolute value where positive values stay positive and negative values become positive.

OpenStudy (anonymous):

from there you would add 3 to both sides and divide it?

OpenStudy (anonymous):

i thought on the negitive one you are suppose to flip it to 3-2x?

OpenStudy (anonymous):

Add 3 and divide will work. With regards to flipping, what do you do with the other side of the equation?